# On the semi-additivity of the $1/2$-symmetric caloric capacity in the plane

### v3

@@ 5.3. Verifying the hypotheses of a $Tb\,$-theorem
which implies, in particular, $\pazocal{P}_{\ast}\nu(\overline{y})>\alpha-\kappa$. If we have had $\overline{y}\in\pazocal{S}_{2}$ we would have obtained the same bound for $\pazocal{P}_{\ast}^{\ast}\nu(\overline{y})$, since $P$ and $P^{\ast}$ share the same C-Z constants. In any case, pick
$\alpha\geq 2\kappa,$
and observe that
@@
The exceptional set $\pazocal{S}$ exhibits additional important properties regarding suppressed kernels (see subsection 3.1), provided their associated 1-Lipschitz function $\Lambda$ satisfies certain conditions. In order to prove them, we present two preliminary results. Fix $\sigma$ any finite Borel measure in $\mathbb{R}^{n+1}$ and $\Lambda:\mathbb{R}^{n+1}\to[0,\infty)$ a 1-Lipschitz function. The following result will be useful in the sequel, and its proof is analogous to that of [Vo, Lemma 8.3].
@@
In light of the previous result, the following lemma also follows and its proof is analogous to that of [To3, Lemma 5.5].
@@
$\big{\rvert}\pazocal{P}_{\Lambda}(\varphi\nu)(\overline{x})\big{\rvert}$ $=\Bigg{\rvert}\sum_{j\in I_{Q}}\frac{\langle T_{0},\varphi_{j}\rangle}{\pazocal{L}^{n+1}(B_{j})}\int_{Q\cap B_{j}}P_{\Lambda}(\overline{x},\overline{y})\varphi(\overline{y})\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})\Bigg{\rvert}$
$\lesssim\sum_{j\in I_{Q}}\frac{1}{r_{j}}\int_{Q\cap B_{j}}\frac{\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})}{|\overline{x}-\overline{y}|^{n}}=:\sum_{j\in I_{Q}}\frac{1}{r_{j}}I_{j}(\overline{x}).$
@@
$\int_{D_{1,j}}\frac{\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})}{|\overline{x}-\overline{y}|^{n}}\lesssim\frac{\pazocal{L}^{n+1}(D_{1,j})}{\text{diam}(B_{j}\cap Q)^{n}}\leq\text{diam}(B_{j}\cap Q)\leq r_{j}.$
@@
$\int_{D_{2,j}}\frac{\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})}{|\overline{x}-\overline{y}|^{n}}\leq\sum_{k\geq 0}\int_{A_{k}}\frac{\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{y})}{|\overline{x}-\overline{y}|^{n}}\lesssim\text{diam}(B_{j}\cap Q)\sum_{k\geq 0}\frac{1}{2^{k}}\leq r_{j}.$
@@
$\int_{Q}\big{\rvert}\pazocal{P}_{\Lambda}(\varphi\nu)(\overline{x})\big{\rvert}\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{x})$ $=\int_{Q}\bigg{\rvert}\int_{F\cap Q}P_{\Lambda}(\overline{x},\overline{y})\varphi(\overline{y})\mathop{}\!\mathrm{d}\nu(\overline{y})\bigg{\rvert}\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{x})$
$\leq\int_{F\cap Q}\bigg{(}\int_{Q}\frac{\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{x})}{|\overline{x}-\overline{y}|^{n}}\bigg{)}|\varphi(\overline{y})|\mathop{}\!\mathrm{d}\mu(\overline{y})\leq\ell(Q)\int_{F\cap Q}|\varphi(\overline{y})|\mathop{}\!\mathrm{d}\mu(\overline{y})$
@@
$\frac{1}{\pazocal{L}^{n+1}(\frac{1}{4}Q)}\int_{\frac{1}{4}Q}\big{\rvert}\pazocal{P}_{\Lambda}(\varphi\nu)(\overline{x})\big{\rvert}\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{x})\lesssim 1,$
@@
    $\mu\big{(}H_{\pazocal{D}(\overline{w})}\cup T_{\pazocal{D}(\overline{w})}\big{)}\leq\delta_{0}\mu(F)$, for some $\delta_{0}\in(0,1)$
@@
($\mathbb{P}\big{(}\big{\{}\overline{w}_{2}\in\Omega\,:\,Q\in\pazocal{D}_{1}\text{ is bad with respect to }\pazocal{D}(\overline{w}_{2})\big{\}}\big{)}\leq\varepsilon_{b}.$
@@
(((([To3, Lemma 5.13])))) Let $\pazocal{D}_{1}=\pazocal{D}(\overline{w}_{1})$ and $\pazocal{D}_{2}=\pazocal{D}(\overline{w}_{2})$ with $\overline{w}_{1},\overline{w}_{2}\in\Omega$. Given $\varepsilon>0$, let $\Lambda:\mathbb{R}^{n+1}\to[\varepsilon,\infty)$ be a 1-Lipschitz function such that
@@
where $\Delta$ is a certain parallelepiped introduced at the beginning of Subsection 5.4. Such expressions were already tackled in [NTVo2] or [HyMar], and they can be deduced from the estimate
@@
where in our setting we may assume $Q$ to be contained in $F\setminus{H}$ and with $M$-thin boundary. This precise bound is covered by assumption 6 in the statement of Theorem 5.17. From this point on, the rest of the proof can be followed as in the remaining sections of [To3, Ch.5] to obtain the desired result. Let us also remark that the proof can be followed almost identically (apart from some dimensional changes) taking into account that we have constructed unique measures $\mu$ and $\nu$ so that assumption 5 holds for the operators $\pazocal{P}$ and $\pazocal{P}^{\ast}$ simultaneously. This enables to obtain relation (5.14), which is essential to carry out the final probabilistic arguments in the proof of the $Tb$-theorem found in [To3, Ch.5] or [Vo].
@@
6. Some general comparability results. The semi-additivity of $\widetilde{\gamma}$ in $\mathbb{R}^{2}$
The main goal of this section is to obtain similar results to Theorem 5.19 but removing assumption $\text{{{A}}}_{\boldsymbol{3}}$. We will be able to do this for any compact set in $\mathbb{R}^{2}$ and for a certain family of sets in $\mathbb{R}^{n+1}$, which in particular contains the usual corner-like Cantor sets presented in [He].
@@ 6.1. The capacity of some parallelepipeds in $\mathbb{R}^{n+1}$
First, let us present some preliminary results that extend that of [MPr, Proposition 6.1]. Let us remark that throughout the forthcoming discussion, any parallelepiped will be closed and will have sides parallel to the coordinate axes.
@@
Let $\mu:=\pazocal{H}^{n}|_{R}$ and pick any $\overline{x}=(x,t)$ with $|t-a|>0$. Observe that
@@
where $u:=y-x$ and $R-x$ denotes a translation of $R$ with respect to the vector $(-x,0)\in\mathbb{R}^{n+1}$. Let us pick $D_{r_{0}}$ an $n$-dimensional ball embedded in $\mathbb{R}^{n}\times\{a\}$, centered at $(0,a)\in\mathbb{R}^{n}\times\{a\}$ and with radius $r_{0}=r_{0}(\overline{x})$ big enough so that $R-x\subset D_{r_{0}}$. Then, there exists a dimensional constant $C>0$ so that
@@
In the last step we have applied [AbS, Eq.15.3.3]. This last expression, thought as a function of $\tau>0$, is bounded uniformly on $(0,+\infty)$ with respect to a dimensional constant (and in fact attains its maximum when $\tau\to 0$). In other words, we deduce that $P_{\text{sy}}\ast\mu(x,t)\lesssim 1$ whenever $|t-a|>0$. On the other hand, it is clear that if $t=a$, then $P_{\text{sy}}\ast\mu(x,t)=0$, meaning that in general
@@
[…]
@@
where $\widetilde{T}_{\varepsilon}\circ\pi$ is smooth and supported on ×$\mathbb{R}\times\Sigma_{\varepsilon}\subset\mathbb{R}\times\mathbb{R}^{n}$, and $C>0$ is some dimensional constant. Therefore, the previous construction implies the following remark:
@@
Let $R\subset\mathbb{R}^{n+1}$ be a parallelepiped and $\ell_{1},\ldots,\ell_{n},\ell_{t}$ its side lengths. Let $R_{\uparrow}$ denote the upper face of $R$ ((((contained in a horizontal affine hyperplane))\,)),
  1. 1))\,))
@@
To prove 1, begin by noticing that $\gamma(R_{\uparrow})=0$ implies $\pazocal{H}^{n}(R_{\uparrow})=0$, by the comments made after Lemma 6.1. Embedding $R_{\uparrow}$ into $\mathbb{R}^{n}$ via $\pi_{x}:\mathbb{R}^{n+1}\to\mathbb{R}^{n}$, $\pi_{x}(\overline{x})=x$, the canonical projection onto the spatial coordinates; we deduce that $Q=\pi_{x}(R)$ is such that $\pazocal{L}^{n}(Q)=0$. But $Q$ is itself a parallelepiped of $\mathbb{R}^{n}$, i.e.
@@
So there must exist some $[a_{j},b_{j}]$ such that $a_{j}=b_{j}$. Hence,
@@
6.2. General comparability for parallelepipeds in $\mathbb{R}^{n+1}$
In light of the above results, we shall obtain a first estimate analogous to Theorem 5.19 for parallelepipeds, where assumption $\text{{{A}}}_{\boldsymbol{3}}$ will not be needed.
@@ 6.2. General comparability for parallelepipeds in $\mathbb{R}^{n+1}$
Let $R\subset\mathbb{R}^{n+1}$ be a closed parallelepiped with sides parallel to the coordinate axes. Then,
@@
Let us denote by $\ell_{1},\ldots,\ell_{n},\ell_{t}$ the side lengths of $R$. Notice that by Lemmas 6.1 and 6.2 we already know the above result if any $\ell_{i}$ is null. So let us assume that
@@
  • Regarding $\text{{{P}}}_{\boldsymbol{1}}$, we get $\frac{5}{8}\pazocal{Q}_{i_{1}}\cap R\neq\varnothing$ for each $i_{1}=1,\ldots,N_{1}$. We wish to study the intersections $2Q_{i_{1}}\cap R$ using the previous property. Notice that $\frac{5}{8}\pazocal{Q}_{i_{1}}\cap R\neq\varnothing$ implies that the sets $2\pazocal{Q}_{i_{1}}\cap R$ are parallelepipeds such that their side with maximal length presents length between $\frac{11}{16}\ell(\pazocal{Q}_{i_{1}})$ and $2\ell(\pazocal{Q}_{i_{1}})$. On the other hand, we distinguish two cases for the sides with minimal length, depending on $\ell_{\text{min}}:=\min\{\ell_{1},\ldots,\ell_{n},\ell_{t}\}$:
@@
    $\ell_{\text{min}}\geq\frac{11}{16}\max_{i_{1}=1,\ldots,N_{1}}\ell(Q_{i_{1}}),$
@@
If 1) occurs we are done, by Theorem 5.19. If 2) occurs, notice that if A) also occurs we get that $R_{1},\ldots R_{N_{1}}$ are parallelepipeds such that all of their respective sides have comparable lengths. In other words, there exists a cube $S_{i_{1}}\subset R_{i_{1}}$ so that $\widetilde{\gamma}(R_{i_{1}})\approx\widetilde{\gamma}(S_{i_{1}})$, for each $i_{1}=1,\ldots,N_{1}$. Then, using that $\text{{{A}}}_{\boldsymbol{3}}$ is not satisfied together with point 2 in Corollary 6.3, we get
@@
Corollary 6.5.
Let $E:=R_{1}\cup R_{2}\cup\cdots\cup R_{N}$ be a finite union of disjoint closed parallelepipeds with sides parallel to the coordinate axes. Then,
@@
The iterative scheme of the proof we have given for Lemma 6.4 can be also applied in this case, but now taking into account the parameter
$\delta:=\min_{i\neq j}\big{\{}\text{dist}(R_{i},R_{j})\big{\}}>0.$

### v4

@@ 5.3. Verifying the hypotheses of a TbTb\,-theorem
which implies, in particular, $\pazocal{P}_{\ast}\nu(\overline{y})>\alpha-\kappa$. If we have had $\overline{y}\in\pazocal{S}_{2}$ we would have obtained the same bound for $\pazocal{P}_{\ast}^{\ast}\nu(\overline{y})$, since $P$ and $P^{\ast}$ share the same CZ constants. In any case, pick $\alpha\geq 2\kappa$ and observe that
@@
The exceptional set $\pazocal{S}$ exhibits additional important properties regarding suppressed kernels (see subsection 3.1), provided their associated 1-Lipschitz function $\Lambda$ satisfies certain conditions. In order to prove them, we present two preliminary results. Fix $\sigma$ any finite Borel measure in $\mathbb{R}^{n+1}$ and $\Lambda:\mathbb{R}^{n+1}\to[0,\infty)$ a 1-Lipschitz function. The next lemma admits an analogous proof to that of [Vo, Lemma 8.3].
@@
The following lemma with an analogous proof to that of [To3, Lemma 5.5].
@@
$\big{\rvert}\pazocal{P}_{\Lambda}(\varphi\nu)(\overline{x})\big{\rvert}$ $=\Bigg{\rvert}\sum_{j\in I_{Q}}\frac{\langle T_{0},\varphi_{j}\rangle}{\pazocal{L}^{n+1}(B_{j})}\int_{Q\cap B_{j}}P_{\Lambda}(\overline{x},\overline{y})\varphi(\overline{y})\mathop{}\!\mathrm{d}\overline{y}\Bigg{\rvert}$
$\lesssim\sum_{j\in I_{Q}}\frac{1}{r_{j}}\int_{Q\cap B_{j}}\frac{\mathop{}\!\mathrm{d}\overline{y}}{|\overline{x}-\overline{y}|^{n}}=:\sum_{j\in I_{Q}}\frac{1}{r_{j}}I_{j}(\overline{x}).$
@@
$\int_{D_{1,j}}\frac{\mathop{}\!\mathrm{d}\overline{y}}{|\overline{x}-\overline{y}|^{n}}\lesssim\frac{\pazocal{L}^{n+1}(D_{1,j})}{\text{diam}(B_{j}\cap Q)^{n}}\leq\text{diam}(B_{j}\cap Q)\leq r_{j}.$
@@
$\int_{D_{2,j}}\frac{\mathop{}\!\mathrm{d}\overline{y}}{|\overline{x}-\overline{y}|^{n}}\leq\sum_{k\geq 0}\int_{A_{k}}\frac{\mathop{}\!\mathrm{d}\overline{y}}{|\overline{x}-\overline{y}|^{n}}\lesssim\text{diam}(B_{j}\cap Q)\sum_{k\geq 0}\frac{1}{2^{k}}\leq r_{j}.$
@@
$\int_{Q}\big{\rvert}\pazocal{P}_{\Lambda}(\varphi\nu)(\overline{x})\big{\rvert}\mathop{}\!\mathrm{d}\overline{x}$ $=\int_{Q}\bigg{\rvert}\int_{F\cap Q}P_{\Lambda}(\overline{x},\overline{y})\varphi(\overline{y})\mathop{}\!\mathrm{d}\nu(\overline{y})\bigg{\rvert}\mathop{}\!\mathrm{d}\overline{x}$
$\leq\int_{F\cap Q}\bigg{(}\int_{Q}\frac{\mathop{}\!\mathrm{d}\overline{x}}{|\overline{x}-\overline{y}|^{n}}\bigg{)}|\varphi(\overline{y})|\mathop{}\!\mathrm{d}\mu(\overline{y})\leq\ell(Q)\int_{F\cap Q}|\varphi(\overline{y})|\mathop{}\!\mathrm{d}\mu(\overline{y})$
@@
$\frac{1}{\pazocal{L}^{n+1}(\frac{1}{4}Q)}\int_{\frac{1}{4}Q}\big{\rvert}\pazocal{P}_{\Lambda}(\varphi\nu)(\overline{x})\big{\rvert}\mathop{}\!\mathrm{d}\overline{x}\lesssim 1,$
@@
    $\mu\big{(}H_{\pazocal{D}(\overline{w})}\cup T_{\pazocal{D}(\overline{w})}\big{)}\leq\delta_{0}\mu(F)$, for some $\delta_{0}\in(0,1)$ .
@@
$\mathbb{P}\big{(}\big{\{}\overline{w}_{2}\in\Omega\,:\,Q\in\pazocal{D}_{1}\text{ is bad with respect to }\pazocal{D}(\overline{w}_{2})\big{\}}\big{)}\leq\varepsilon_{b}.$
@@
(([To3, Lemma 5.13])) Let $\pazocal{D}_{1}=\pazocal{D}(\overline{w}_{1})$ and $\pazocal{D}_{2}=\pazocal{D}(\overline{w}_{2})$ with $\overline{w}_{1},\overline{w}_{2}\in\Omega$. Given $\varepsilon>0$, let $\Lambda:\mathbb{R}^{n+1}\to[\varepsilon,\infty)$ be a 1-Lipschitz function such that
@@
where $\Delta$ is a certain parallelepiped introduced at the beginning of Subsection 5.4. Such expressions were already tackled in [NTVo2] and [HyMar], and they can be deduced from the estimate
@@
where in our setting we may assume $Q$ to be contained in $F\setminus{H}$ and with $M$-thin boundary. This precise bound is covered by assumption 6 in the statement of Theorem 5.17. From this point on, the rest of the argument can be followed as in the remaining sections of [To3, Ch.5]. Let us also remark that the proof can be followed almost identically (apart from some dimensional changes) taking into account that we have constructed unique measures $\mu$ and $\nu$ so that assumption 5 holds for the operators $\pazocal{P}$ and $\pazocal{P}^{\ast}$ simultaneously. This enables to obtain relation (5.14), which is essential to carry out the final probabilistic arguments in the proof of the $Tb$-theorem found in [To3, Ch.5] or [Vo].
@@
6. The general comparability in $\mathbb{R}^{n+1}$
The main goal of this section is to obtain similar results to Theorem 5.19 but removing assumption $\text{{{A}}}_{\boldsymbol{3}}$. We will be able to do this for any compact set in $\mathbb{R}^{n+1}$.
@@ 6.1. The capacity of some parallelepipeds in $\mathbb{R}^{n+1}$
First, let us present some preliminary results that extend that of [MPr, Proposition 6.1]. Let us remark that throughout the forthcoming discussion, any parallelepiped will have sides parallel to the coordinate axes.
@@
Let $\mu:=\pazocal{H}^{n}|_{R}$ and pick any $\overline{x}=(x,t)$ with |ta|>0|t-a|>0. Observe that
@@
where u:=yxu:=y-x and RxR-x denotes a translation of $R$ with respect to the vector $(-x,0)\in\mathbb{R}^{n+1}$. Let us pick $D_{r_{0}}$ an $n$-dimensional ball embedded in $\mathbb{R}^{n}\times\{a\}$, centered at $(0,a)\in\mathbb{R}^{n}\times\{a\}$ and with radius $r_{0}=r_{0}(\overline{x})$ big enough so that $R-x\subset D_{r_{0}}$. Then, there exists a dimensional constant $C>0$ so that
@@
In the last step we have applied [AbS, Eq.15.3.3]. This last expression, thought as a function of $\tau>0$, is bounded uniformly on $(0,+\infty)$ with respect to a dimensional constant (and in fact attains its maximum when $\tau\to 0$). In other words, we deduce that $P_{\text{sy}}\ast\mu(x,t)\lesssim 1$ whenever |ta|>0|t-a|>0. On the other hand, it is clear that if $t=a$, then $P_{\text{sy}}\ast\mu(x,t)=0$, meaning that in general
@@
$\pi:$ $\quad\mathbb{R}^{n+1}\longrightarrow\mathbb{R}^{n},\hskip 56.9055pt\iota:\mathbb{R}^{n}\longrightarrow\hskip 7.68236pt\mathbb{R}^{n+1},$
$(x_{1},x)\;\hskip 1.42271pt\longmapsto x\hskip 83.93553ptx\hskip 6.25963pt\longmapsto(0,x)$
@@
where $\widetilde{T}_{\varepsilon}\circ\pi$ is smooth and supported on $\mathbb{R}\times\Sigma_{\varepsilon}\subset\mathbb{R}\times\mathbb{R}^{n}$, and $C>0$ is some dimensional constant. Therefore, the previous construction implies the following remark:
@@
Let $R\subset\mathbb{R}^{n+1}$ be a parallelepiped and $\ell_{1},\ldots,\ell_{n},\ell_{t}$ its side lengths. Let $R_{\uparrow}$ denote the upper face of $R$ ((contained in a horizontal affine hyperplane)\,),
  1. 1)\,)
@@
To prove 1, begin by noticing that $\gamma(R_{\uparrow})=0$ implies $\pazocal{H}^{n}(R_{\uparrow})=0$, by the comments made after Lemma 6.1. Embedding $R_{\uparrow}$ into $\mathbb{R}^{n}$ via $\pi_{x}:\mathbb{R}^{n+1}\to\mathbb{R}^{n}$, $\pi_{x}(\overline{x})=x$, the canonical projection onto the spatial coordinates; we deduce that $Q=\pi_{x}(R)$ is such that $\pazocal{L}^{n}(Q)=0$. But $Q$ is itself a parallelepiped of $\mathbb{R}^{n}$ (that we assume to be closed without loss of generality), i.e.
@@
So there must exist some $a_{j},b_{j}]$ such that $a_{j}=b_{j}$. Hence,
@@
6.2. General comparability in $\mathbb{R}^{n+1}$
In light of the previous results, we first obtain an estimate analogous to Theorem 5.19 for parallelepipeds, without requiring assumption $\text{{{A}}}_{\boldsymbol{3}}$. Then, we establish a similar result for finite unions of closed cubes of the same size with disjoint interiors. Since $\text{{{A}}}_{\boldsymbol{2}}$ can be assumed without loss of generality, the proof of the main theorem follows.
@@ 6.2. General comparability in $\mathbb{R}^{n+1}$
Let $R\subset\mathbb{R}^{n+1}$ be a parallelepiped with sides parallel to the coordinate axes. Then,
@@
Let us denote by $\ell_{1},\ldots,\ell_{n},\ell_{t}$ the side lengths of $R$. Observe that for our purposes, we may assume $R$ to be closed without loss of generality. Notice that by Lemmas 6.1 and 6.2 we already know the above result if any $\ell_{i}$ is null. So let us assume that
@@
  • Regarding $\text{{{P}}}_{\boldsymbol{1}}$, we get $\frac{5}{8}\pazocal{Q}_{i_{1}}\cap R\neq\varnothing$ for each $i_{1}=1,\ldots,N_{1}$. We wish to study the intersections $2\pazocal{Q}_{i_{1}}\cap R$ using the previous property. Notice that $\frac{5}{8}\pazocal{Q}_{i_{1}}\cap R\neq\varnothing$ implies that the sets $2\pazocal{Q}_{i_{1}}\cap R$ are parallelepipeds such that their side with maximal length presents length between $\frac{11}{16}\ell(\pazocal{Q}_{i_{1}})$ and $2\ell(\pazocal{Q}_{i_{1}})$. On the other hand, we distinguish two cases for the sides with minimal length, depending on $\ell_{\text{min}}:=\min\{\ell_{1},\ldots,\ell_{n},\ell_{t}\}$:
@@
    $\ell_{\text{min}}\geq\frac{11}{16}\max_{i_{1}=1,\ldots,N_{1}}\ell(\pazocal{Q}_{i_{1}}),$
@@
If 1) occurs we are done, by Theorem 5.19. If 2) occurs, notice that if A) also occurs we get that $R_{1},\ldots R_{N_{1}}$ are parallelepipeds such that all of their respective sides have comparable lengths. In other words, there exists a cube $P_{i_{1}}\subset R_{i_{1}}$ so that $\widetilde{\gamma}(R_{i_{1}})\approx\widetilde{\gamma}(P_{i_{1}})$, for each $i_{1}=1,\ldots,N_{1}$. Then, using that $\text{{{A}}}_{\boldsymbol{3}}$ is not satisfied together with point 2 in Corollary 6.3, we get
@@
Lemma 6.5.
Let $E:=Q_{1}\cup Q_{2}\cup\cdots\cup Q_{N}$ be a finite union of closed cubes of equal size with disjoint interiors and sides parallel to the coordinate axes. Then,
@@
First, we observe that it is enough to prove the result if $Q_{1},\ldots,Q_{N}$ are disjoint. Indeed, if they were not, write for each $j=1,\ldots,N$,
$Q_{j}:=\prod_{i=1,\ldots,n,t}[a_{ij},b_{ij}]\quad\text{and}\quad m_{ij}:=\frac{b_{ij}-a_{ij}}{2},\quad i=1,\ldots,n,t,$
and for each $0<\lambda<1$ define the contracted concentric cube
$Q_{j}^{\lambda}:=\prod_{i=1,\ldots,n,t}[a_{ij}+\lambda m_{ij},b_{ij}-\lambda m_{ij}],$
as well as $E^{\lambda}:=Q_{1}^{\lambda}\cup Q_{2}^{\lambda}\cup\cdots\cup Q_{N}^{\lambda}$. Observe that the family $\{Q_{1}^{\lambda},\dots,Q_{N}^{\lambda}\}$ consists of closed disjoint cubes of the same size, and therefore $\widetilde{\gamma}(E^{\lambda})\lesssim\widetilde{\gamma}_{+}(E^{\lambda})$. Moreover, by construction we get that for any $0<\lambda_{1}\leq\lambda_{2}<1$, $E^{\lambda_{2}}\subset E^{\lambda_{1}}$. Then, by monotonicity
$\sup_{0<\lambda<1}\widetilde{\gamma}(E^{\lambda})=\lim_{\lambda\to 0}\widetilde{\gamma}(E^{\lambda})=\widetilde{\gamma}(\mathring{E}),$
where $\mathring{E}$ denotes the interior of $E$. We have a analogous identity for $\widetilde{\gamma}_{+}$. Since for our particular choice of $E$ we have $\widetilde{\gamma}_{+}(\mathring{E})\approx\widetilde{\gamma}_{+}(E)$ (see Remark 6.3), we are done:
$\widetilde{\gamma}_{+}(E)\lesssim\widetilde{\gamma}_{+}(\mathring{E})\leq\widetilde{\gamma}(\mathring{E})=\sup_{0<\lambda<1}\widetilde{\gamma}(E^{\lambda})\lesssim\sup_{0<\lambda<1}\widetilde{\gamma}_{+}(E^{\lambda})\leq\widetilde{\gamma}_{+}(E).$
So we shall assume that $Q_{1},\ldots,Q_{N}$ are disjoint. In this context, the iterative scheme of the proof given for Lemma 6.4 can be also applied, but now taking into account the parameter
$\delta:=\min_{i\neq j}\big{\{}\text{dist}(Q_{i},Q_{j})\big{\}}>0.$